OpenStudy (anonymous):

A long distance telephone company advertised the rates of 5 minutes for 85¢ and 10 minutes for $1.10. Use a linear equation to find the cost of a 25 minute phone call.

OpenStudy (anonymous):

find the equation of the line through the points (5,85) and (10,110)

OpenStudy (anonymous):

slope is \[\frac{110-85}{10-5}=\frac{25}{5}=5\]

OpenStudy (anonymous):

use point slope formula to get \[y-y_1=m(x-x_1)\] \[y-110=5(x-10)\] \[y-110=5x-50\] \[y=5x+60\]

OpenStudy (anonymous):

interpretation is you get charged 5 cents a minute and 60 cents to connect the call. replace x by 25 to get \[y=5\times 25+60=125+60=185\] or $1.85